Andrew is two years older that Bianca if their ages have a sum of 28 years, how old is each one?
Found 2 solutions by Alan3354, macston:
Answer by Alan3354(69443)

Andrew is two years older that Bianca if their ages have a sum of 28 years, how old is each one?
-----------
A + B = 28
A - B = 2
----------------- Add

Answer by macston(5194)

.
A=Andrew's age; B=Bianca's age
.
A=B+2yrs . Use this to substitute for A below
.
A+B=28 yrs . Substitute for A from above.
.
(B+2yrs)+B=28yrs . Subtract 2 yrs from each side.
.
2B=26 yrs . Divide each side by 2.
.
B=13 yrs
.
ANSWER 1: Bianca is 13 years old.
.
A=B+2yrs=13yrs+2yrs=15yrs
.
ANSWER 2: Andrew is 15 years old.
.
CHECK:
.
A+B=28yrs
.
15yrs+13yrs=28yrs
.
28yrs=28yrs
.